Session: Nov 2001 (session 3)
Wavelength: 5.0 cm


To give feedback on this experiment, click here

Experiment Feedback

\(Other\) (A. Cattani at Fri Nov 9 18:26:04 MET 2001) Minor failures

Medicina (A. Cattani at Fri Nov 9 18:26:44 MET 2001) Minor failures

Torun (K.Borkowski at Fri Nov 9 19:58:59 MET 2001) Failure
Our Penny&Giles recorder upgraded to two heads and FS 9.5.2 (and earlier versions) continue refuse to cooperate.

Effelsberg (A. Kraus at Sun Nov 11 13:47:02 MET 2001) Minor failures
Wind and snowfall prohibited antenna movement between UT1700-1900 and UT0325-0400. Due to problems with the drive program, scans between UT2234-2305, UT1044-1101, UT1308-1310.

Onsala (John Conway at Mon Nov 12 09:48:39 MET 2001) Success
No known problems

Jodrell_Bank (Alastair Gunn at Wed Nov 14 10:20:11 MET 2001) Minor failures
Telescope azimuth drive failed at 00:31, restarted at 00:41, missed scan on DA193. Tsys measurements in VC04 not working due to component failure. No other known problems.

Cambridge (Alastair Gunn at Wed Nov 14 10:21:11 MET 2001) Minor failures
FS crashed at 00:06, restarted on DA193 scan at 00:37. Vacuum failed at 07:10, restarted on 3C84 at 07:22. Vacuum failed again soon after 11:00. Discovered tape had snapped at about 8000 ft. Shortage of tapes meant rest of schedule was discarded. Broken tape sent to correlator.